Ammonia, NH3, is a weak base with a Kb value of 1.8×10−5.

What is the pH of a 0.320 mol L−1 ammonia solution?

Answer is 11.38

Part B

What is the percent ionization of ammonia at this concentration?Express your answer with the appropriate units.
Answer #1

Since Ammonia is a weak base, in water, it can't ionise 100%. Thus it will form a conjugate acid and OH- ion.

  NH3 + H2O <----> NH4+ + OH-

Here Kb= the equilibrium constant of the base=  [NH4+][OH-]/[NH3]

where [NH4+]= concentration of NH4+ at eqilibrium

[OH-] = concentration of OH- at eqilibrium

[NH3]=  concentration of NH3 at eqilibrium

If we form the ICE table, then NH3 + H20 ----------------> NH3+ + OH-

I= 0.320 mol L−1 0 0

C= -x +x +x

E= 0.320-x x x

Thus Kb= x*x/o.320-x

or 1.8×10−5 = x2 / o.320-x

or x= 0.023 mol L-1

Here x= the amount of NH3 reduced from initial= amount of ammonia ionised qt equillibrium.

Then the % of ionisation= amount of ammonia ionised qt equillibrium/ amount of NH3 taken initially * 100

=  0.023 mol L-1/ 0.320 mol L−1 *100

= 7.18 %